# Your Paper Title Here

### v1

@@ 3 Experiments with Simulated Data
Our simulation study benchmarks all debiasing methods and monitors the upstream and downstream analyses while maintaining the firewall between them. The simulation uses a canonical directed acyclic graph (DAG), shown in Figure 4. It is canonical in the sense that the EO-ML literature tends to assume this structure (Sakamoto et al., 2025). It illustrates the various research phases and their corresponding data availability.
In phase 0, the data-generation process (approximating reality) unfolds in the direction of the arrows. In the upstream phase, only the blue-nodes data are available, with finite samples of $Y_{i}$, to produce $\hat{Y_{i}}$ for the entire population $N$. In the downstream, only red nodes are available: $\hat{Y_{i}}$ is now available for $N$, jointly with $C_{i}$ and $A_{is}$. They enable inference on the estimand $\tau_{s}$. (We suppress the index $s$ henceforth.) Appendix IV delineates our simulation further.
Figure 4: A panel illustrating data availability in the two phases.
@@
We focus on two key objectives: minimizing error in the upstream task and minimizing causal bias in the downstream task. To assess error, we measure how closely the predicted values match the true values of $\tau$, using Mean Absolute Error (MAE). A lower MAE indicates a more accurate estimation of the outcomes, and thus, a more accurate downstream ATE estimate. To assess bias, we examine the extent to which the upstream outcome prediction errors vary systematically when changing the magnitude of $\tau$. This examination entails regressing the predicted $\hat{\tau}$ values on the true $\tau$ values and analyzing the calibration slope. If the estimates are unbiased, the slope should be close to one. Deviations from one indicate systematic over-/underestimation across the range of $\tau$.
We can formally test deviations from 111 with an $F$-test: $F$-statistics from the above regression quantify the strength of the relationship between the predicted treatment effects and the true treatment effects. These $F$-statistic results (Table 1) indicate that Tweedie’s correction is the only method that yields ATE estimates for which we cannot reject the null hypothesis that these estimates are unbiased. Figure A.III.1 visualizes these slope results.
Method MAE Slope $F$-statistic $p$-value
Tweedie’s 0.04 0.995 1.22 0.303
LCC 0.05 1.008 5.75 0.006
PPI (10%) 0.19 0.985 6.441 0.003
Ratledge 0.37 0.641 32.93 0.000
Naive 0.48 0.535 6420.43 0.000
Table 1: Simulation performance metrics. $F$-statistics and $p$-values are the statistical test for evaluating the null hypothesis that the slope value is 1. Failure to reject indicates that the slope estimate intervals contain the true ATE.
4 Experiments with Real Data
In this section, we conduct two types of downstream-real experiments. First, we will evaluate how well our debiasing methods succeed in descriptive tasks, especially those involving group aggregation. Second, we will apply our procedure to causal tasks. These two tasks cover the main use cases of our EO-ML-driven poverty data.
@@ 4 Experiments with Real Data
Here, $Y_{i}$ denotes the International Wealth Index (IWI), a continuous asset-based wealth score (0–100) aggregated at the cluster level to mitigate noise (Burke et al., 2021; Zhu et al., 2025). The covariate array $\mathbf{X}_{i}$ consists of Landsat satellite imagery (30m resolution, six multispectral bands) centered on each DHS cluster (6.72 km × 6.72 km tiles), processed by taking the three-year median over cloud-free pixels leading up to the survey date.
Two models were fitted for each fold, one using MSE and one using the loss introduced by Ratledge et al. (2022). The models shared the same ResNet-50 architecture and were initialized with weights pretrained on the SSL4EO-L dataset (Stewart et al., 2023). For more details about model training, see Appendix II. The resulting MSE models all have an $R^{2}$ score between 0.74 and 0.76 on their hold-out fold, and an average slope of 0.79 . Their combined predictions can be seen in Figure 2. The Ratledge models have a slope of 0.99, but at the cost of reduced $R^{2}$ scores between 0.67 and 0.70.
@@
As expected, the MSE of the naive model exhibits shrinkage behavior: it tends to overpredict the wealth of poorer regions and underpredict the wealth of richer regions. As shown in Figure 5, Tweedie’s correction accurately compensates for this, not only resulting in less biased estimates (regression slope of 0.83 to 0.90), but also decreases the MAE from 2.67 to 2.39. In fact, if we consider only regions with at least 100 clusters, the Tweedie-corrected estimates exhibit virtually no shrinkage bias, with a slope of 0.99. Appendix III delineates this analysis.
Figure 5: Mean residuals and Tweedie’s corrections for each ADM2 region in the dataset. regions that are overestimated in the left plot (blue) are shown to be correctly adjusted down in the right plot (red), and vice versa.
@@
To assess the practical utility of our correction methods in a setting where the true treatment effect is unknown, we construct an evaluation framework based on real-world development projects. Specifically, we leverage the same dataset of geo-referenced interventions funded by the World Bank and the Chinese government as was used in Malik et al. (2021); Conlin (2024). Each project is tagged with an associated aid sector $s\in\mathcal{S}$ (e.g., Health, Water Supply and Sanitation, or Women in Development), which defines a downstream trial $D_{\text{Trial }s}=\{(\bm{X}_{i},\widehat{Y}_{i},A_{is}):i\in\mathcal{I}_{\textrm{Trial }s}\}$. Here, $A_{is}\in\{0,1\}$ indicates exposure to a $s$ intervention at location $i$, $\bm{X}_{i}$ are satellite features, and $\widehat{Y}_{i}$ are IWI predictions from the upstream-trained model. Figure A.III.2 depicts the spatial distribution of treatments across the continent.
Since the treatment effects of these interventions are unknown, we deem this analysis a test of external validity by comparing EO-ML predicted wealth values with observed ones. Our goal is to assess whether our bias-corrected predictions can reflect causal estimates similar to those from observed data.
@@
For each funder-sector combination, we estimate the average treatment effect as the difference in mean outcomes between treated/control villages. We compute this using both observed IWI values and the model-predicted (and correction-adjusted) IWI values.
@@
See Figure 6 for results. Each point is a causal study. While the $x$-axis shows what would have happened when using the label $Y_{i}$ in these studies, the $y$-axis shows the corresponding causal effect using $\hat{Y_{i}}$ with its corresponding adjustment approach. Orange lines check the correlation among these data. A score of one is perfect. The analysis shows that Tweedie’s correction (panel d) comes close to this desiderata, with a correlation of 0.9980.9980.998; and it does so with the lowest MAE of 0.570.570.57. Had we used the naive predictions (panel a), the correlation would be 0.9580.9580.958 and the MAE would be 0.690.690.69. Assuming the PPI set up, with 10% fresh data available, we would have achieved a 0.9810.9810.981; and it does so with the lowest MAE of 0.730.730.73. The PPI performance is, however, a function of available data in the downstream, and panel f shows that with increasing fresh data, the PPI correction will outperform our Tweedie. But PPI is fresh-data hungry, and thus, this approach would come at a cost of deploying more costly surveys in Africa—costing downstream teams millions of USD de Pieuchon et al. (2025).

### v2

@@ 2. Tweedie’s correction
Our simulation study benchmarks all debiasing methods and monitors the upstream and downstream analyses while maintaining the firewall between them. The simulation uses a canonical directed acyclic graph (DAG), shown in Figure 3. It is canonical in the sense that the EO-ML literature tends to assume this structure (sakamoto2024scoping), while illustrating the various research phases and corresponding data needs.
In phase 0, the data-generation process (approximating reality) unfolds in the direction of the arrows. In the upstream phase, only the blue-nodes data are available, with finite samples of $Y_{i}$, to produce $\hat{Y_{i}}$ for the entire population NN. In the downstream, only red nodes are available: $\hat{Y_{i}}$ is now available for NN, jointly with $C_{i}$ and $A_{is}$. They enable inference on the estimand $\tau_{s}$. (We suppress the index $s$ henceforth.) Appendix IV delineates our simulation further.
Figure 3: A panel illustrating the problem setup.
@@
We focus on two key objectives in the downstream task: minimizing error and minimizing causal bias. To assess error, we measure how closely the predicted values match the true values of $\tau$, using Mean Absolute Error (MAE). A lower MAE indicates a more accurate recovery of the true treatment effects. To assess bias, we examine the calibration slope from regressing the predicted effects $\hat{\tau}$ on the true $\tau$ values and analyzing the calibration slope. If the estimates are unbiased, the slope should be close to one. Intuitively, a slope of one means predicted and true scales align (no global shrinkage) so downstream treatment effects are not attenuated purely by scaling. Deviations from one indicate systematic over-/underestimation across the range of $\tau$. We report both the slope and its standard error, checking whether the 95% confidence interval includes one. Figure A.III.1 visualizes these slope results.
Results (Table 1) show that Tweedie’s correction achieves the lowest MAE and yields slope estimates whose confidence interval contains one, indicating no detectable attenuation bias. In contrast, both the Ratledge and naive approaches yield calibration slopes significantly below one, indicating systematic underestimation of effects.
Method MAE Slope $\pm$ SE 1 $\in$ 95% CI?
Tweedie’s 0.04 0.995 $\pm$ 0.006 $\checkmark$
LCC 0.05 1.008 $\pm$ 0.007 $\checkmark$
PPI (10%) 0.19 0.985 $\pm$ 0.028 $\checkmark$
Ratledge 0.37 0.641 $\pm$ 0.024 $\times$
Naive 0.48 0.535 $\pm$ 0.004 $\times$
Table 1: Performance on simulated treatment-effect estimation. Tweedie’s correction yields the lowest error and an unbiased slope; Ratledge and the naive approach show substantial attenuation bias.
Experiments with Real Data
In this section, we conduct two types of downstream-real experiments. First, we will evaluate how well our debiasing methods succeed in descriptive tasks, especially those involving group aggregation. Second, we will apply our procedure to causal tasks. These two tasks cover the main use cases of our EO-ML-driven poverty data (see Figure 5).
@@
Here, $Y_{i}$ denotes the International Wealth Index (IWI), a continuous asset-based wealth score (0–100) aggregated at the cluster level to mitigate noise (burkeUsingSatelliteImagery2021; pmlr-v275-zhu25a). The covariate array $\mathbf{X}_{i}$ consists of Landsat satellite imagery (30 m resolution, six multispectral bands) centered on each DHS cluster (6.72 km × 6.72 km tiles), processed by taking the three-year median over cloud-free pixels leading up to the survey date.
Two models were fitted for each fold, one using MSE and one using the loss introduced by ratledge_using_2022. The models shared the same ResNet-50 architecture and were initialized with weights pretrained on the SSL4EO-L dataset (stewart2023ssl4eoldatasetsfoundationmodels). For more details about model training, see Appendix II. The resulting MSE models all have an $R^{2}$ score between 0.74 and 0.76 on their hold-out fold, and an average slope of 0.79 . Their combined predictions can be seen in Figure 2. The Ratledge models have a slope of 0.99, but at the cost of reduced $R^{2}$ scores between 0.67 and 0.70.
@@
As expected, the MSE of the naive model exhibits shrinkage behavior: it tends to overpredict the wealth of poorer regions and underpredict the wealth of richer regions. As shown in Figure 4, Tweedie’s correction accurately compensates for this, not only resulting in less biased estimates (regression slope of 0.83 to 0.90), but also decreases the MAE from 2.67 to 2.39. In fact, if we consider only regions with at least 100 clusters, the Tweedie-corrected estimates exhibit virtually no shrinkage bias, with a slope of 0.99. Appendix III explicates.
Figure 4: Mean residuals (Left) and adjustments from Tweedie’s method (Right) for each ADM2 region in the dataset. Regions that are overestimated in the left plot (blue) are shown to be correctly adjusted down in the right plot (red), and vice versa.
@@
To assess the practical utility of our correction methods in a setting where the true treatment effect is unknown, we construct an evaluation framework based on real-world development projects. Specifically, we leverage the same dataset of geo-referenced interventions funded by the World Bank and the Chinese government as was used in (malik2021banking; Conlin2024). Each project is tagged with an associated aid sector $s\in\mathcal{S}$ (e.g., Health, Water Supply and Sanitation, or Women in Development), which defines a downstream trial $D_{\text{Trial }s}=\{(\mathbf{X}_{i},\hat{Y}_{i},A_{is}):i\in\mathcal{I}_{\textrm{Trial }s}\}$. Here, $A_{is}\in\{0,1\}$ indicates exposure to a sector-$s$ intervention at location $i$, $\mathbf{X}_{i}$ are satellite features, and $\hat{Y}_{i}$ are IWI predictions from the upstream-trained model. Figure A.III.2 depicts the spatial distribution of treatments across the continent.
Since the treatment effects of these interventions are unknown, we deem this analysis a test of external validity by comparing ML predicted wealth values with observed ones. Our goal is to assess whether bias-corrected predictions can reflect causal estimates similar to those from groundtruth.
@@
For each funder-sector combination, we estimate the average treatment effect as the difference in mean outcomes between treated/control villages, for simplicity. We compute this using both observed IWI values and the model-predicted (and correction-adjusted) IWI values. The former provides a benchmark estimate using labeled data, while the latter evaluates the debiasing methods applied to predictions.
@@
See Figure 5 for results. Each point is a causal study. While the $x$-axis shows what would have happened when using the label $Y_{i}$ in these studies, the $y$-axis shows the corresponding causal effect using $\hat{Y_{i}}$ with its corresponding adjustment approach. Orange lines check the correlation among these data. A score of one is perfect. The analysis shows that Tweedie’s correction (panel d) comes close to this desiderata, with a correlation of $0.998$; and it does so with the lowest MAE of $0.57$. Had we used the naive predictions (panel a), the correlation would be $0.958$ and the MAE would be $0.69$. Assuming the PPI set up, with 10% fresh data available, we would have achieved a $0.981$; and it does so with the lowest MAE of $0.73$. The PPI performance is, however, a function of available data in the downstream, and panel f shows that with increasing fresh data, the PPI correction will outperform our Tweedie. However, since PPI demands substantial amounts of such data, this method would necessitate costly new surveys, resulting in large expenditures for downstream teams (depieuchon2025benchmarkingdebiasingmethodsllmbased).